# Multi-Type Context-Aware Conversational Recommender Systems via Mixture-of-Experts

### v1

@@ 1 Introduction
One of the challenges is the heterogeneity of different external data. For instance, knowledge graphs are structured data while conversations and item reviews are unstructured data. There is a semantic gap among multi-type and heterogeneous data. Furthermore, different external data often belong to different semantic spaces and thus are difficult to align. One possible solution to fuse different external data is contrastive learning. However, contrastive learning requires the same entries contained in all different external data to calculate the contrastive loss. This is not always fulfilled.
In this work, we propose a Multi-type Context-aware Conversational Recommender System, called MCCRS, to effectively fuse multi-type contextual information via mixture-of-experts so as to improve conversational recommender systems. MCCRS incorporates both structured information and unstructured information, including the structured knowledge graph, unstructured conversation history, and unstructured item reviews. We carefully design our model as a mixture-of-experts structure, which consists of a ChairBot and several experts. Each expert is specialized in a particular domain, i.e., one external source. The ChairBot coordinates multiple experts and generates the final results. Compared with previous approaches, the advantages of MCCRS are at least four-fold: First, our MCCRS model relieves the challenge of data heterogeneity and semantic fusion, as well as relaxes the limitation of contrastive learning. Second, our MCCRS model takes advantage of different contextual information and the specialization of different experts followed by a ChairBot breaks the model bottleneck on a single contextual information. Third, with each expert specialized in a particular domain, it is more easily traceable and analyzes who is responsible when the model generates a good or bad result. Fourth, our MCCRS model is more easily extendable as additional experts can be added easily to incorporate more additional external sources.
In summary, the main contributions of this work are threefold [^1] We will make our code publicly available upon acceptance:
@@ 2 Related Work
The other category, known as dialog-based conversational recommender system [35, 7, 36, 10, 37, 38, 39], is based on human-generated dialogs. Unlike anchor-based conversational recommender systems simulating conversations based on extracted anchor text, dialog-based conversational recommender systems concentrate on generating human-like responses while making accurate recommendations. In the early stages, as there are no publicly available datasets for dialog-based conversational recommender systems, Li et al. [40] propose a conversational recommender system dataset in the movie domain, named ReDial. Subsequently, Hayati et al. [41] propose the INSPIRED dataset tailored for conversational recommender systems with sociable recommendation strategies. Dialogues in these datasets typically encompass items and entities that can be linked to knowledge bases like DBpedia. Consequently, many existing studies incorporate external knowledge graphs to augment conversation information and accurately capture user preference [8, 3, 4, 12, 42, 43, 44, 45]. These studies typically employ graph neural networks to encode the knowledge graph and user preferences, alongside a dialog management module to guide the conversation. For example, Chen et al. [3] introduce the entity-oriented knowledge graph (DBpedia [46]), to discern user intentions. They employ Relational Graph Convolutional Networks (R-GCNs) to learn entity representations from knowledge graphs [47] for item recommendations and utilize the Transformer framework to generate natural language responses. Based on the work of Chen et al. [3], which mainly focuses on the entity-oriented knowledge graph, Zhou et al. [4] expand upon this by introducing an additional knowledge graph, i.e., the word-oriented knowledge graph – ConceptNet [48]. They propose a novel knowledge graph-based semantic fusion model for conversational recommender systems, harnessing the potential of both knowledge graphs. Subsequently, Zhou et al. [14] propose a contrastive learning approach[49] from a coarse-to-fine perspective to enhance data semantic fusion. In order to mitigate redundant information in knowledge graphs, Sarkar et al. [5] introduce different subgraphs to enhance recommendation performance for conversational recommender systems. To make use of multiple reasoning paths in knowledge graphs, Ma et al. [6] and Zhou et al. [50] propose conversational recommendation models to perform multi-hop reasoning in knowledge graphs to track shifts in users’ interests. To alleviate the limitation of incomplete knowledge graphs, Zhang et al. [51] present a variational reasoning approach [15] with dynamic knowledge reasoning over incomplete knowledge graphs. Beyond knowledge graphs, pre-trained language models are also deployed to enhance conversational recommender systems [52, 9, 53, 54, 55, 56, 57, 58]. Wang et al. [9] finetune the large-scale pre-trained language models and propose a unified pre-trained language model-based framework to deal with the low-resource challenge in conversational recommender systems. Similarly, Wang et al. [52] apply a pre-trained language model and propose a conversational recommendation model on top of the prompt learning paradigm. Dao et al. [57] introduce a demonstration-enhanced prompt learning method for conversational recommendation by retrieving and learning from demonstrations based on dialogue contexts. Another promising direction in conversational recommendation involves the integration of side information [13, 11]. For example, Lu et al. [13] incorporate the item reviews to improve the recommendation performance in conversational recommender systems. Li et al. [59] leverage multi-aspect user information to learn multi-aspect user preferences, while Yang et al. [11] introduce item meta information to improve item representations.
@@ 3.2 Conversation Expert
Afterward, we would combine their calculation with a residual link. The reason for the residual connection is that the network is more difficult to train as it becomes deeper, and thus employing such a technique along with layer normalization and dropouts assists with the training.
@@
where sub-layer is either PFFN or MultiHead in Equation 2.
@@
where $j$ is the conversation index, and $y_{ij}$ indicates the ground-truth label of items. After the training, we collect both the predicted item probability $P(i)$ and hidden embedding $\mathbf{h}^{N}$ for the ChairBot.
@@ 3.3 Graph Expert
Given it is difficult to comprehensively understand user preferences based solely on conversational context, the inclusion of external knowledge is necessary to encode user preferences. For dialogs in conversational recommender systems, item mentions and item-related entities can be extracted to construct external knowledge graphs. Inspired by the previous studies [3, 4], we introduce a knowledge graph sourced from DBpedia [63] to encode structural and relational information in the knowledge graph. Specifically, we perform entity linking [64] to map the item mentions and item-related entities in the dataset to DBpedia. With the help of the external knowledge graph, it enables us to model user preferences more accurately.
A knowledge graph $\mathcal{G}$ (i.e., DBpedia) comprises an entity set $\mathcal{E}$ and a relation set $\mathcal{R}$. The knowledge graph $\mathcal{G}$ stores semantic facts as triples $<e_{1},r,e_{2}>$, where $e_{1},e_{2}\in\mathcal{E}$ represents items or item-related entities and $r\in\mathcal{R}$ represents the relation between $e_{1}$ and $e_{2}$. For example, as shown in Figure 2, for a target movie $e_{i^{*}}$, it might be associated with entity $e_{1}$ with one type of relationship, while $e_{2}$ and $e_{3}$ are of another, such as $e_{1}$ being the genre while $e_{2}$ and $e_{3}$ are the actors in the movies.
@@
where $\mathbf{n}_{i}$ represents the learned embedding for item $i$. Again, a cross-entropy loss is used for the graph expert similar to the conversation expert (i.e., Equation 5). Finally, the graph expert produces the item probability $P(i)$ and node embeddings for the ChairBot.
@@ 3.4 Review Expert
Also, a cross-entropy loss is used for optimizing the review expert similar to the conversation expert (i.e., Equation 5). Again, the item probability $P(i)$ and review embeddings are passed to the ChairBot.
@@ 3.5 Mixture-of-Experts Recommender
where ${b={C,G,R}}$, and MLP is a linear layer.
@@
where $P_{C}(i)$, $P_{G}(i)$, and $P_{R}(i)$ are item probabilities generated by conversations expert, graph expert, and review expert, respectively. To fine-tune the entire item recommender, we also apply a cross-entropy loss in the Mixture-of-Experts recommender module (i.e., Equation 5) to fine-tune the representations and improve the performance. [^3] One can also use the joint loss for all experts and ChairBot to optimize the recommender. We decided to train all experts separately because we observed better performance in the experiments when training them separately.
@@ 4.1.3 Implementation Details
Table 3: Recommendation results on ReDial. ‘*’ indicates significant improvements over the best baseline (Fisher random test, $p$-value $<0.05$). The best performances are highlighted in bold, while the second-best are underlined (Unless otherwise reported, we use ‘*’ to indicate significant improvements, bold values to indicate the best performances, and underlined values to indicate the second-best performances throughout the paper). Our proposed MCCRS method significantly outperforms SOTA baselines on the recommendation task on ReDial.
@@ 4.2.1 Performance on Recommendation Task
For the recommendation performances on ReDial in Table 3, we observe that our proposed MCCRS significantly outperforms all the baselines on all metrics on the ReDial dataset. MCCRS outperforms the best classical recommendation baseline BERT by 307%, 114%, and 148% in terms of Recall@1, Recall@10, and Recall@50, respectively. Compared with the SOTA conversational recommendation baselines $\text{C}^{2}$-CRS and VRICR, MCCRS outperforms $\text{C}^{2}$-CRS by 7.5%, 7.3%, and 16.2% while outperforms VRICR by 5.6%, 2.5%, and 16.5%, in terms of Recall@1, Recall@10, and Recall@50, respectively. Similarly, observe from Table 4 for the recommendation performances on INSPIRED, our proposed model MCCRS still significantly outperforms all the baselines on all metrics. MCCRS achieves a significant improvement of 136.4% on Recall@1, 53.6% on Recall@10, and 51.5% on Recall@50 over the best classical baseline BERT. Compared with the SOTA conversational recommendation baseline $\text{C}^{2}$-CRS, MCCRS achieves a significant improvement of 15.6% on Recall@1, 13.6% on Recall@10, and 24.6% on Recall@50. The above observations indicate the effectiveness of our MCCRS model by incorporating multiple experts. MCCRS utilizes multi-type external information, including conversational history, knowledge graph, and reviews, which is helpful in understanding the conversation context. The effective Mixture-of-Experts framework to fuse the multi-type external information enhances the data representations and improves the recommendation performance. For conversational baselines, we find that KBRD, and KGSF perform better than ReDial, due to their use of external knowledge graphs, which aid in interpreting user intentions more effectively. RevCore achieves better recommendation performance than KBRD and KGSF. This might be because RevCore incorporates additional review information on items. $\text{C}^{2}$-CRS outperforms RevCore, as it performs semantic fusing techniques with more external information.
@@ 4.2.2 Performance on Conversation Task
Table 7 shows the performance comparison of human evaluation on response generation. From the results, we find that ReDial outperforms the typical Transformer by incorporating a pre-trained RNN encoder. $\text{C}^{2}$-CRS performs the best among baselines on the reported human evaluation metrics. This might be because that $\text{C}^{2}$-CRS facilitates semantic fusion of data by using contrastive learning, which improves the responses.
@@ 4.3 Ablation Study (RQ2)
We explore the contributions of different components within our model in this section. We perform an ablation study on the ReDial dataset, by considering six variants of the model, including (1) “w/o conversation” removes the conversation expert from the framework; (2) “w/o graph” removes the graph expert; (3) “w/o review” removes the review expert; (4) “conversation only” only involves the single conversation expert; (5) “graph only” only involves the single graph expert; (6) “review only” only involves the single review expert. Observe from Table 8, among the three experts, “graph only” achieves the highest performance than the other two single experts. Also, we find that removing any of the experts, either conversation expert, graph expert, or review expert, results in lower performance. This indicates that all three experts contributed to the final performance. After removing the graph expert, the performance drops more than the other two experts. This indicates the graph expert contributes the most among the three experts, highlighting the importance of the graph expert. This might be because the graph expert benefits from the structure of the knowledge graph. Moreover, although we observe that “review only” and “conversation only” achieve relatively low performance, removing the conversation expert or review expert results in a decrease in performance. This suggests that adding conversations or reviews will marginally improve the recommendation accuracy. The reviews, although did not play a major role, do help with the recommendation task even if reviews are missing for some items.
@@ 4.4 Parameter Sensitivity (RQ3)
In this section, we explore how our proposed model is affected by its main parameters, including the mask proportion and hidden dimensionality.
@@ 4.4.2 Hidden Dimensionality
Table 10 illustrates the effect of embedding size on model performance. Overall, the recommendation performance generally declines as dimensionality increases from 32 to 256. This is probably due to overfitting. The proposed model achieves its best performance with an embedding size of 32.
@@
Table 10: Effect of embedding size.
Embedding Size Recall@1 Recall@10 Recall@50
@@ 5 Conclusion and Future Work
One limitation of this work is that we use entities from the DBpedia knowledge graph following previous work of conversational recommender system [3, 5, 4, 6], the extracted entities may not be 100% accurate. Therefore, it is valuable to explore the conversational recommender system by integrating and modeling the uncertainty and noise in conversation contexts. Also, we do not incorporate sentiment analysis of items, entities, or reviews in the multi-type contextual data. In future work, we plan to explore conversational recommender systems by integrating and modeling the sentiment (e.g., positive or negative) from the conversation and reviews in future work.

### v2

@@ 1 Introduction
One of the challenges is the heterogeneity of different external data. For instance, knowledge graphs are structured data while conversations and item reviews are unstructured data. There is a semantic gap between multi-type and heterogeneous data. Furthermore, different external data often belong to different semantic spaces and thus are difficult to align. One possible solution to fuse different external data is contrastive learning [14, 16]. However, contrastive learning requires the same entries contained in all different external data to calculate the contrastive loss. This is not always fulfilled.
In this work, we propose a Multi-type Context-aware Conversational Recommender System, called MCCRS, to effectively fuse multi-type contextual information via mixture-of-experts so as to improve conversational recommender systems. MCCRS incorporates both structured information and unstructured information, including the structured knowledge graph, unstructured conversation history, and unstructured item reviews. We carefully design our model as a mixture-of-experts structure, which consists of a ChairBot and several experts. Each expert specializes in a particular domain, i.e., one external source. The ChairBot coordinates multiple experts and generates the final results. Compared with previous approaches, the advantages of MCCRS are at least four-fold: First, our MCCRS model relieves the challenge of data heterogeneity and semantic fusion, as well as relaxes the limitation of contrastive learning. Second, our MCCRS model takes advantage of different contextual information and the specialization of different experts followed by a ChairBot breaks the model bottleneck on a single contextual information. Third, with each expert specialized in a particular domain, it is more easily traceable and analyzes who is responsible when the model generates a good or bad result. Fourth, our MCCRS model is more easily extendable as additional experts can be added easily to incorporate more additional external sources.
In summary, the main contributions of this work are threefold:
@@ 2 Related Work
The other category, known as dialog-based conversational recommender system [36, 37, 10, 38, 39, 40], is based on human-generated dialogs. Unlike anchor-based conversational recommender systems simulating conversations based on extracted anchor text, dialog-based conversational recommender systems concentrate on generating human-like responses while making accurate recommendations. In the early stages, as there are no publicly available datasets for dialog-based conversational recommender systems, Li et al. [41] propose a conversational recommender system dataset in the movie domain, named ReDial. Subsequently, Hayati et al. [42] propose the INSPIRED dataset tailored for conversational recommender systems with sociable recommendation strategies. Dialogues in these datasets typically encompass items and entities that can be linked to knowledge bases like DBpedia. Consequently, many existing studies incorporate external knowledge graphs to augment conversation information and accurately capture user preference [8, 3, 4, 12, 43, 44, 45, 46]. These studies typically employ graph neural networks to encode the knowledge graph and user preferences, alongside a dialog management module to guide the conversation. For example, Chen et al. [3] introduce the entity-oriented knowledge graph (DBpedia [47]), to discern user intentions. They employ Relational Graph Convolutional Networks (R-GCNs) to learn entity representations from knowledge graphs [48] for item recommendations and utilize the Transformer framework to generate natural language responses. Based on the work of Chen et al. [3], which mainly focuses on the entity-oriented knowledge graph, Zhou et al. [4] expand upon this by introducing an additional knowledge graph, i.e., the word-oriented knowledge graph – ConceptNet [49]. They propose a novel knowledge graph-based semantic fusion model for conversational recommender systems, harnessing the potential of both knowledge graphs. Subsequently, Zhou et al. [14] propose a contrastive learning approach[16] from a coarse-to-fine perspective to enhance data semantic fusion. In order to mitigate redundant information in knowledge graphs, Sarkar et al. [5] introduce different subgraphs to enhance recommendation performance for conversational recommender systems. To make use of multiple reasoning paths in knowledge graphs, Ma et al. [6] and Zhou et al. [50] propose conversational recommendation models to perform multi-hop reasoning in knowledge graphs to track shifts in users’ interests. To alleviate the limitation of incomplete knowledge graphs, Zhang et al. [51] present a variational reasoning approach [15] with dynamic knowledge reasoning over incomplete knowledge graphs. Beyond knowledge graphs, pre-trained language models are also deployed to enhance conversational recommender systems [52, 9, 53, 54, 55, 56, 57, 58, 59, 60]. Wang et al. [9] finetune the large-scale pre-trained language models and propose a unified pre-trained language model-based framework to deal with the low-resource challenge in conversational recommender systems. Similarly, Wang et al. [52] apply a pre-trained language model and propose a conversational recommendation model on top of the prompt learning paradigm. Dao et al. [57] introduce a demonstration-enhanced prompt learning method for conversational recommendation by retrieving and learning from demonstrations based on dialogue contexts. Wei et al. [60] propose a novel multi-modal semantic graph prompt learning framework for conversational recommender systems by considering collaborative information with multi-modal features. An et al. [59] propose a novel model of contextual disentanglement for conversational recommender systems by disentangling focus and background information from complex conversation contexts. Another promising direction in conversational recommendation involves the integration of side information [13, 11]. For example, Lu et al. [13] incorporate the item reviews to improve the recommendation performance in conversational recommender systems. Li et al. [61] leverage multi-aspect user information to learn multi-aspect user preferences, while Yang et al. [11] introduce item meta information to improve item representations.
@@ 3.2 Conversation Expert
Afterward, we would combine their calculation with a residual link. The reason for the residual connection is that the network is more difficult to train as it becomes deeper, and thus employing such a technique, along with layer normalization and dropout, assists with the training.
@@
where the sublayer is either PFFN or MultiHead in Equation 2.
@@
where $j$ is the conversation index, and $y_{ij}$ indicates the ground-truth label of items. After the training, we collect both the predicted item probability $P_{C}(i)$ and hidden embedding $\mathbf{h}^{N}$ for the ChairBot.
@@ 3.3 Graph Expert
Given that it is difficult to comprehensively understand user preferences based solely on conversational context, the inclusion of external knowledge is necessary to encode user preferences. For dialogs in conversational recommender systems, item mentions and item-related entities can be extracted to construct external knowledge graphs. Inspired by the previous studies [3, 4], we introduce a knowledge graph sourced from DBpedia [65] to encode structural and relational information in the knowledge graph. Specifically, we perform entity linking [66] to map the item mentions and item-related entities in the dataset to DBpedia. With the help of the external knowledge graph, it enables us to model user preferences more accurately.
A knowledge graph $\mathcal{G}$ (i.e., DBpedia) comprises an entity set $\mathcal{E}$ and a relation set $\mathcal{R}$. The knowledge graph $\mathcal{G}$ stores semantic facts as triples $e_{1},r,e_{2}>$, where $e_{1},e_{2}\in\mathcal{E}$ represents items or item-related entities and $r\in\mathcal{R}$ represents the relation between $e_{1}$ and $e_{2}$. For example, as shown in Figure 2, for a target movie $e_{i^{*}}$, it might be associated with entity $e_{1}$ with one type of relationship, while $e_{2}$ and $e_{3}$ are of another, such as $e_{1}$ being the genre while $e_{2}$ and $e_{3}$ are the actors in the movies.
@@
where $\mathbf{n}_{i}$ represents the learned embedding for item $i$. Again, a cross-entropy loss is used for the graph expert similar to the conversation expert (i.e., Equation 5). Finally, the graph expert produces the item probability $P_{G}(i)$ and node embeddings for the ChairBot.
@@ 3.4 Review Expert
Also, a cross-entropy loss is used for optimizing the review expert similar to the conversation expert (i.e., Equation 5). Again, the item probability $P_{R}(i)$ and review embeddings are passed to the ChairBot.
@@ 3.5 Mixture-of-Experts Recommender
where b=C,G,R{b={C,G,R}}, and MLP is a linear layer.
@@
where $P_{C}(i)$, $P_{G}(i)$, and $P_{R}(i)$ are item probabilities generated by the conversation expert, graph expert, and review expert, respectively. To fine-tune the entire item recommender, we also apply a cross-entropy loss in the Mixture-of-Experts recommender module (i.e., Equation 5) to fine-tune the representations and improve the performance. [^2] One can also use the joint loss for all experts and ChairBot to optimize the recommender. We decided to train all experts separately because we observed better performance in the experiments when training them separately.
@@ 4.1.3 Implementation Details
Table 3: Recommendation results on ReDial. ‘*’ indicates significant improvements over the best baseline (Fisher random test, $p$-value <0.05<0.05). The best performances are highlighted in bold, while the second-best are underlined (Unless otherwise reported, we use ‘*’ to indicate significant improvements, bold values to indicate the best performances, and underlined values to indicate the second-best performances throughout the paper). Our proposed MCCRS method significantly outperforms SOTA baselines on the recommendation task on ReDial.
@@ 4.2.1 Performance on Recommendation Task
For the recommendation performances on ReDial in Table 3, we observe that our proposed MCCRS significantly outperforms all the baselines on all metrics on the ReDial dataset. MCCRS outperforms the best classical recommendation baseline BERT by 307%, 114%, and 148% in terms of Recall@1, Recall@10, and Recall@50, respectively. Compared with the SOTA conversational recommendation baselines $\text{C}^{2}$-CRS and VRICR, MCCRS outperforms $\text{C}^{2}$-CRS by 7.5%, 7.3%, and 16.2% while outperforms VRICR by 5.6%, 2.5%, and 16.5%, in terms of Recall@1, Recall@10, and Recall@50, respectively. Similarly, observe from Table 4 for the recommendation performances on INSPIRED, our proposed model, MCCRS, still significantly outperforms all the baselines on all metrics. MCCRS achieves a significant improvement of 136.4% on Recall@1, 53.6% on Recall@10, and 51.5% on Recall@50 over the best classical baseline BERT. Compared with the SOTA conversational recommendation baseline $\text{C}^{2}$-CRS, MCCRS achieves a significant improvement of 15.6% on Recall@1, 13.6% on Recall@10, and 24.6% on Recall@50. The above observations indicate the effectiveness of our MCCRS model by incorporating multiple experts. MCCRS utilizes multi-type external information, including conversational history, knowledge graph, and reviews, which is helpful in understanding the conversation context. The effective Mixture-of-Experts framework to fuse the multi-type external information enhances the data representations and improves the recommendation performance. For conversational baselines, we find that KBRD, and KGSF perform better than ReDial, due to their use of external knowledge graphs, which aid in interpreting user intentions more effectively. RevCore achieves better recommendation performance than KBRD and KGSF. This might be because RevCore incorporates additional review information on items. $\text{C}^{2}$-CRS outperforms RevCore, as it performs semantic fusing techniques with more external information.
@@ 4.2.2 Performance on Conversation Task
Table 7 shows the performance comparison of human evaluation on response generation. From the results, we find that ReDial outperforms the typical Transformer by incorporating a pre-trained RNN encoder. $\text{C}^{2}$-CRS performs the best among baselines on the reported human evaluation metrics. This might be because $\text{C}^{2}$-CRS facilitates semantic fusion of data by using contrastive learning, which improves the responses.
@@ 4.3 Ablation Study (RQ2)
We explore the contributions of different components within our model in this section. We perform an ablation study on the ReDial dataset, by considering six variants of the model, including (1) “w/o conversation” removes the conversation expert from the framework; (2) “w/o graph” removes the graph expert; (3) “w/o review” removes the review expert; (4) “conversation only” only involves the single conversation expert; (5) “graph only” only involves the single graph expert; (6) “review only” only involves the single review expert. Observe from Table 8, among the three experts, “graph only” achieves the highest performance than the other two single experts. Also, we find that removing any of the experts, either the conversation expert, graph expert, or review expert, results in lower performance. This indicates that all three experts contributed to the final performance. After removing the graph expert, the performance drops more than the other two experts. This indicates the graph expert contributes the most among the three experts, highlighting the importance of the graph expert. This might be because the graph expert benefits from the structure of the knowledge graph. Moreover, although we observe that “review only” and “conversation only” achieve relatively low performance, removing the conversation expert or review expert results in a decrease in performance. This suggests that adding conversations or reviews will marginally improve the recommendation accuracy. The reviews, although did not play a major role, do help with the recommendation task even if reviews are missing for some items.
@@ 4.4 Parameter Sensitivity (RQ3)
In this section, we explore how our proposed model is affected by its main parameters, including the mask proportion and hidden dimensionality, on the ReDial dataset.
@@ 4.4.2 Hidden Dimensionality
Table 10 illustrates the effect of hidden dimensionality (embedding size) on model performance. Overall, the recommendation performance generally declines as dimensionality increases from 32 to 256. This is probably due to overfitting. The proposed model achieves its best performance with a hidden dimension of 32.
@@
Table 10: Effect of hidden dimensionality.
Hidden Dimensionality Recall@1 Recall@10 Recall@50
@@ 5 Conclusion and Future Work
One limitation of this work is that we use entities from the DBpedia knowledge graph following previous work of conversational recommender system [3, 5, 4, 6], the extracted entities may not be 100% accurate. Therefore, it is valuable to explore the conversational recommender system by integrating and modeling the uncertainty and noise in conversation contexts. Also, we do not incorporate sentiment analysis of items, entities, or reviews in the multi-type contextual data. In future work, we plan to explore conversational recommender systems by integrating and modeling the sentiment information (e.g., positive or negative) from the conversation and reviews in future work.
@@
Last, while MCCRS achieves strong performance in benchmark datasets and effectively fuses multi-type contextual information, its mixture-of-experts structure may lead to increased computational complexity, which may pose challenges in real-world deployment scenarios when involving large-scale or real-time recommendation tasks. In future work, we plan to explore potential techniques, such as knowledge distillation for model compression [72] and expert pruning, for improving model efficiency and scalability. Also, for the mixture-of-experts structure, we combine the results of multiple experts with a followed ChairBot. One can also use other combination strategies of multiple experts, e.g., attention-based gating mechanisms, or a hierarchical structure that sequentially refines expert outputs.
6 Acknowledgement
We would like to thank Qingtian Cao for her contribution to the early exploration and discussion of this work. This research was supported by the National Natural Science Foundation of China (62402093) and, the Sichuan Science and Technology Program (2025ZNSFSC0479). This work was also supported in part by the National Natural Science Foundation of China under grants U20B2063 and 62220106008, and the Sichuan Science and Technology Program under Grant 2024NSFTD0034.